OpenStudy (anonymous):

2x+1=square root of (3x^2+5x+3) how do I solve this?

OpenStudy (anonymous):



OpenStudy (anonymous):

lmfao, that's the answer?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how'd you get it?

OpenStudy (anonymous):

Magic :) haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

It's just simplifying

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

\[(2x+1)^2=3x^2+5x+3, \text{where } 3x^2+5x+3 \ge 0\]

OpenStudy (anonymous):

Then you simplify the quadratic, for 2 solutions of x. Check them both against the inequality to see which, or neither, or both work. :)

OpenStudy (anonymous):

so there can be two answers?

OpenStudy (anonymous):

-1/3 is not the answer btw

OpenStudy (anonymous):

okay, then once you get your two answers you check it, then what?

OpenStudy (anonymous):

potentially if both solutions make 3x2+5x+3 greater than or equal to 0

OpenStudy (anonymous):

An answer that statisfies both the equality and the inequality is THE answer.

OpenStudy (anonymous):

okay, so if you check it and it works once you plug it back in, its the answer?

OpenStudy (anonymous):

You should check BOTH solutions of the equality against the inequality. (this example does have 2 solutions)

OpenStudy (anonymous):

oh okay, i have another one similar to this next question. do u mind helping me with that too

OpenStudy (anonymous):

Sure do you want me to verify your solution to this one first?

OpenStudy (anonymous):

I already put in the wrong one. lol. I'm too trusting. but i'm going to do it correctly next

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the next one is x+2=square root of (-1-2x)

OpenStudy (anonymous):

I got down to two answers -1 and -5 but i checked and -5 doesnt work. so it's -1?

OpenStudy (anonymous):

So we have\[x+2=\sqrt{-1-2x}\] we square both sides and keep in mind that:\[-1-2x \ge 0\] \[(x+2)^2=-1-2x\]\[x^2+4x+4=-1-2x\]\[x^2+6x+5=0\]

OpenStudy (anonymous):

yeah, that's what i got when I simplified

OpenStudy (anonymous):

You got it right -1

OpenStudy (amoodarya):

I think x=2 is right answer not x=-1 check it !

OpenStudy (amoodarya):